# ReQISC: A Reconfigurable Quantum Computer Microarchitecture and Compiler Co-Design

### v1

@@ 4. The ReQISC microarchitecture
While any 2Q gate within the Weyl chamber can theoretically be realized through Hamiltonian steering (zhang2005generation; lin2022let; mckinney2023parallel), pulse concatenation or optimization (song2019generation; omran2019generation; gokhale2020optimized), these methods often require complex control sequences or fail to achieve time-optimal gate implementation. In contrast, we propose a unified, straightforward, and unequivocally superior control scheme that directly implements (up to local equivalence) arbitrary 2Q gates in optimal time, as depicted by Figure 3 and Algorithm 1. Inspired by the AshN gate scheme (chen2024one) which was limited to $XX+YY$ coupling, our scheme accommodates arbitrary coupling Hamiltonians for any quantum processor, whether it be superconducting, trapped-ion, or other platforms, significantly broadening applicability. Time optimality (hammerer2002characterization) refers to the theoretical lower bound of evolution time to achieve an equivalent 2Q unitary given the coupling Hamiltonian, with tunable parallel local drives. This is essential for minimizing the qubits’ exposure to decoherence and also for reducing overall runtime.
@@
Output : 2Q gate time $\tau$; Local Hamiltonians $H_{1},H_{2}\in\mathcal{H}({C^{2\times 2}})$; 1Q gate corrections $A_{1},A_{2},B_{1},B_{2}\in\mathbf{SU}(2)$
1
2 $(x,y,z,\,A^{\prime}_{1},A^{\prime}_{2},\,B^{\prime}_{1},B^{\prime}_{2})\leftarrow\textsc{CanonicalDecompose}(U)$;
[…]
4
[…]
6
[…]
8 $\tau\leftarrow\min\{\tau_{1},\tau_{2}\}$;
9
10if $\tau_{2}<\tau_{1}$ then
11 $x\leftarrow\frac{\pi}{2}-x$; $z\leftarrow-z$; $\tau_{0},\,\tau_{+},\,\tau_{-}\leftarrow\tau_{0}^{\prime},\,\tau_{+}^{\prime},\,\tau_{-}^{\prime}$;
12
13 end if
14
15if $\tau=\tau_{0}$ then
16 $S_{1}\leftarrow\mathrm{sinc}^{-1}(\frac{\sin(y-z)}{(b-c)\tau})/\tau$ where $S_{1}\geq b-c$;
17 $S_{2}\leftarrow\mathrm{sinc}^{-1}(\frac{\sin(y+z)}{(b+c)\tau})/\tau$ where $S_{2}\geq b+c$;
18 $\Omega_{1}\leftarrow\frac{1}{2}\sqrt{S_{1}^{2}-(b-c)^{2}};\,\Omega_{2}\leftarrow\frac{1}{2}\sqrt{S_{2}^{2}-(b+c)^{2}};\,\delta\leftarrow 0$;
19
20else
21 $\mathrm{lhs}(\alpha,\beta,\eta,t)\coloneqq\frac{(1-\alpha)(1+\alpha+\beta)-(1-\alpha+\beta)\eta}{(1-\alpha+\beta)(1-\eta+\alpha+2\beta)}e^{-i(2+2\beta-\eta)t}+\frac{\beta(1+\alpha+\beta)-(1-\alpha+\beta)\eta}{(1-\alpha+\beta)(2\alpha+\beta-\eta)}e^{i(\eta-2\alpha)t}+\frac{(1+\alpha-\eta)\eta-\beta(1-\alpha-\eta)}{(2\alpha+\beta-\eta)(1+\alpha+2\beta-\eta)}e^{i(2\alpha+2\beta-\eta)t}$;
22 $\mathrm{rhs}(x,y,z)\coloneqq e^{i(x-y-z)}-e^{i(y-x-z)}+e^{i(z-x-y)}$;
23 if $\tau=\tau_{+}$ then
24    $x^{\prime},\,y^{\prime},\,z^{\prime}\leftarrow x+c\tau,\,y+c\tau,\,c\tau-z$;
25    $\eta\leftarrow(a-b)/(a+c);\quad t\leftarrow(a+c)\tau$;
26
27    $(\alpha,\beta)\leftarrow\textsc{NSolve}(\mathrm{lhs}(\alpha,\beta,\eta,t)=\mathrm{rhs}(x^{\prime},y^{\prime},z^{\prime}))$ where $(\alpha,\beta)\in[0,1]\times[0,\infty)$ and $\alpha+\beta\geq\eta$;
28    $\Omega_{1}\leftarrow 0$; $\Omega_{2}\leftarrow(a+c)\sqrt{(1-\alpha)\beta(1-\eta+\alpha+\beta)}$;
29    $\delta\leftarrow-(a+c)\sqrt{\alpha(1+\beta)(\alpha+\beta-\eta)}$;
30
31 else
32    $x^{\prime},\,y^{\prime},\,z^{\prime}\leftarrow x-c\tau,\,y-c\tau,\,z-c\tau$;
33    $\eta\leftarrow(a-b)/(a-c);\quad t\leftarrow(a-c)\tau$;
34    $(\alpha,\beta)\leftarrow\textsc{NSolve}(\mathrm{lhs}(\alpha,\beta,\eta,t)=\mathrm{rhs}(x^{\prime},y^{\prime},z^{\prime}))$ where $(\alpha,\beta)\in[0,1]\times[0,\infty)$ and $\alpha+\beta\geq\eta$;
35    $\Omega_{1}\leftarrow(a-c)\sqrt{(1-\alpha)\beta(1-\eta+\alpha+\beta)}$; $\Omega_{2}\leftarrow 0$;
36    $\delta=(a-c)\sqrt{\alpha(1+\beta)(\alpha+\beta-\eta)}$;
37
38   end if
39
40 end if
[…]
@@ 36    $\delta=(a-c)\sqrt{\alpha(1+\beta)(\alpha+\beta-\eta)}$;
42 $H_{1},\,H_{2}\leftarrow U_{1}H^{\prime\prime}_{1}U_{1}^{\dagger}-H^{\prime}_{1},\,U_{2}H^{\prime\prime}_{2}U_{2}^{\dagger}-H^{\prime}_{2}$;
43 $A_{1},\,A_{2}\leftarrow A^{\prime}_{1}(A^{\prime\prime}_{1})^{\dagger}U_{1}^{\dagger},A^{\prime}_{2}(A^{\prime\prime}_{2})^{\dagger}U_{2}^{\dagger}$;
@@ 4.1. Protocol design (Hamiltonian description)
To illustrate the general applicability of our scheme, we consider the most general 2Q coupling Hamiltonian and local Rabi drives (zhang2005generation) for a pair of resonant qubits ($\omega_{q_{1}}=\omega_{q_{2}}=\omega$), with experimentally tunable pulse parameters:
@@
where $(a,b,c)$ are the canonical coupling coefficients (bennett2002optimal) with $a\geq b\geq\lvert c\rvert$. We define the coupling strength $g:=a+b+\lvert c\rvert$. For flux-tunable superconducting qubits, in the drive frame after the rotating wave approximation, $a=b=g/2$, $c=0$, $A_{i}$ is the amplitude of a square wave envelope for a sinusoidal drive and $\delta_{i}:=\frac{1}{2}(\omega_{d_{i}}-\omega)$ is the drive detuning. We actually show that it is sufficient to set the drive frequencies to be the same, so $\delta_{1}=\delta_{2}=\delta$. For other physical platforms, $H_{d_{1}}$ and $H_{d_{2}}$ may be realized via other means (e.g., lasers in Rydberg atoms or trapped ions). In this case, $A_{i}$ and $\delta_{i}$ may correspond to other physical parameters. Our framework can also be easily adapted to other possible $H_{d_{1}},H_{d_{2}}$ (for example $YI$ and $IY$). By defining the symmetric and anti-symmetric combinations of drive amplitudes (for mathematical convenience) $\Omega_{1,\,2}=-\frac{1}{4}{(A_{1}\pm A_{2})}$, we can rewrite the local drives as
(2) $H_{1}=(\Omega_{1}+\Omega_{2})XI+\delta\,ZI,\,H_{2}=(\Omega_{1}-\Omega_{2})IX+\delta\,IZ.$
@@
Algorithm 1 describes the details of our gate scheme. A target gate $U$ and the device 2Q Hamiltonian $H$ are provided as inputs. The goal is to determine the local Hamiltonians $H_{1}$ and $H_{2}$, with the interaction duration $\tau$, such that $e^{-i\tau(H+H_{1}+H_{2})}\sim U$. Specifically, the Weyl coordinates $(x,y,z)$ of $U$ are obtained via canonical decomposition (line 1), while the interaction Hamiltonian $H$, if not already in its native normalized form, is transformed into its normal form (dur2001entanglement) to obtain its canonical coefficients $(a,b,c)$ (line 1). The parameters $(x,y,z)$ and $(a,b,c)$ are then fed to the subsequent solving procedure that computes $H_{1},H_{2},\tau$. With additional 1Q gate corrections $A_{1,\,2}$ and $B_{1,\,2}$ (line 1 and 1), $(A_{1}\otimes A_{2})e^{-i\tau(H+H_{1}\otimes I+I\otimes H_{2})}(B_{1}\otimes B_{2})$ exactly implements $U$.
Herein we give some intuition for Algorithm 1 while leaving the detailed proof in Appendix A. First, the interaction time $\tau$ is determined as the maximum value among $\{\tau_{0},\tau_{+},\tau_{-}\}$. Which of the three is largest dictates the subscheme used: (1) $\tau_{0}$, that is, no-detuning (ND, $\delta=0$); (2) $\tau_{+}$ for equal-amplitudes with opposite signs (EA+, $\Omega_{1}=0$); (3) $\tau_{-}$ for equal-amplitudes with same sign (EA-, $\Omega_{2}=0$). The three subschemes correspond to three distinct regions that partition the Weyl chamber (chen2024one). This partitioning of the Weyl chamber is convenient in terms of physical implementation but not unique. Each subscheme [^1] Note that $\mathrm{sinc}(x)$ (line 1 and 1) is not invertible (bijective), so $S_{1}$ can correspond to any of the preimages. However, in hardware implementations, we typically select the smallest possible value to minimize the required pulse amplitude. Similarly, we opt for the smallest possible values when solving the equations in lines 1 and 1. involves transcendental equations to calculate amplitudes and/or detuning, which lacks analytical solutions and must be solved numerically. We can also prove Algorithm 1 provides a time-optimal (hammerer2002characterization; dur2001entanglement) realization for any 2Q gates in Appendix A.
Figure 4. Hardware implementation of the ReQISC microarchitecture. (a) 2Q gate mirroring. The near-identity corner with a $\mathcal{L}^{1}$ coordinate norm bounded by $r$ is mirrored to the corner near $\mathrm{SWAP}$. (b) Gate time landscape, computed under the $\mathrm{XY}$ coupling Hamiltonian $H=\frac{g}{2}(XX+YY)$, where the circle size indicates the corresponding gate duration within the Weyl chamber. (c-d): Pulse control subschemes, under $\mathrm{XY}$ and $\mathrm{XX}$ couplings, respectively. (e) Local drive amplitudes required for representative gate families. The scaling factor $s$ represents the specific equivalent canonical gate within the gate family, e.g., $\mathrm{B}^{s}\sim\mathrm{Can}(s\frac{\pi}{4},s\frac{\pi}{8},0)$. $\mathrm{iSWAP}$ family requires no local drives so it is not shown in the figure; $\mathrm{CNOT}$ family and $\mathrm{B}$ family require only one-side drive ($A_{1}\neq 0$); $\mathrm{SWAP}$ family requires both-side drives ($A_{1},A_{2}\neq 0$). Frequency-related quantities ($A_{1}$, $A_{2}$, $\delta$) are normalized by the coupling strength $g$.

### v2

@@ 4. The ReQISC microarchitecture
While any 2Q gate can theoretically be realized through Hamiltonian steering (Zhang and Whaley, 2005; Lin et al., 2022; McKinney et al., 2023), pulse concatenation or optimization (Song et al., 2019; Omran et al., 2019; Gokhale et al., 2020), they often require complex control sequences or fail to achieve time-optimality. In contrast, we propose a unified and simple control scheme that directly implements (up to local equivalence) arbitrary 2Q gates in optimal time, as depicted by Figure 3 and Algorithm 1. Inspired by the AshN scheme (Chen et al., 2024), which was limited to $\mathrm{XY}$ coupling, our scheme accommodates arbitrary coupling Hamiltonians for any quantum processor, whether it be superconducting, trapped-ion, or other platforms, significantly broadening applicability. Time optimality (Hammerer et al., 2002) refers to the theoretical lower bound of evolution time required to implement a locally equivalent 2Q unitary under a specific coupling Hamiltonian and tunable local drives. This is essential for mitigating decoherence and reducing overall runtime.
From a systems perspective, the hardware provides a fixed “datapath” determined by its physical coupling Hamiltonian (e.g., constant $\mathrm{XX}+\mathrm{YY}$ or $\mathrm{XX}$ interaction). ReQISC acts as the control logic that configures the local drives to steer this datapath. The generation of these control signals follows a rigorous logic flow, as formalized in Figure 3(b). It consists of three stages: ① Instruction decoding: The target gate $U$ is decomposed into its canonical coordinates $(x,y,z)$, mapping it to a point in the geometric instruction space (Weyl chamber). ② Exception handling: We identify “near-identity” $\mathrm{SU}(4)$ gates that physically require arbitrarily large energies to execute optimally due to control singularities. Instead of executing them directly, the compiler transforms these gates into their “mirrored” version by appending a logical SWAP, thereby avoiding the singularity. The effects of extra $\mathrm{SWAP}$ gates can be resolved by just tracking qubit mapping changes. ③ Micro-op synthesis: Based on the instruction’s coordinates, the logic selects one of three optimal execution modes—no-detuning (ND), equal-amplitude+ (EA+), and equal-amplitude- (EA-). This selection is analogous to selecting the optimal datapath configuration in a CPU to minimize latency (gate duration $\tau$).
@@
Output : 2Q gate time $\tau$; Local drive Hamiltonians $H_{1},H_{2}\in\mathcal{H}({C^{2\times 2}})$; 1Q gate corrections $A_{1},A_{2},B_{1},B_{2}\in\mathbf{SU}(2)$
$(x,y,z,\,A^{\prime}_{1},A^{\prime}_{2},\,B^{\prime}_{1},B^{\prime}_{2})\leftarrow\textsc{CanonicalDecompose}(U)$;
$(a,b,c,\,H^{\prime}_{1},H^{\prime}_{2},\,U_{1},U_{2})\leftarrow\textsc{NormalForm}(H)$;
$\tau_{0},\,\tau_{+},\,\tau_{-}\leftarrow\frac{x}{a},\,\frac{x+y-z}{a+b-c},\,\frac{x+y+z}{a+b+c}$;
$\tau_{1}\leftarrow\max\{\tau_{0},\tau_{+},\tau_{-}\}$;
$\tau_{0}^{\prime},\,\tau_{+}^{\prime},\,\tau_{-}^{\prime}\leftarrow\frac{\frac{\pi}{2}-x}{a},\,\frac{\frac{\pi}{2}-x+y+z}{a+b-c},\,\frac{\frac{\pi}{2}-x+y-z}{a+b+c}$;
$\tau_{2}\leftarrow\max\{\tau_{0}^{\prime},\tau_{+}^{\prime},\tau_{-}^{\prime}\}$;
$\tau\leftarrow\min\{\tau_{1},\tau_{2}\}$;
if $\tau_{2}<\tau_{1}$ then
 $x\leftarrow\frac{\pi}{2}-x$; $z\leftarrow-z$;
 $\tau_{0},\,\tau_{+},\,\tau_{-}\leftarrow\tau_{0}^{\prime},\,\tau_{+}^{\prime},\,\tau_{-}^{\prime}$;
end if
if $\tau=\tau_{0}$ then
 $S_{1}\leftarrow\mathrm{sinc}^{-1}\bigl(\frac{\sin(y-z)}{(b-c)\tau}\bigr)/\tau$ where $S_{1}\geq b-c$;
 $S_{2}\leftarrow\mathrm{sinc}^{-1}\bigl(\frac{\sin(y+z)}{(b+c)\tau}\bigr)/\tau$ where $S_{2}\geq b+c$;
 $\Omega_{1}\leftarrow\frac{1}{2}\sqrt{S_{1}^{2}-(b-c)^{2}};\,\Omega_{2}\leftarrow\frac{1}{2}\sqrt{S_{2}^{2}-(b+c)^{2}}\;\delta\leftarrow 0$;
else
 $\mathrm{lhs}(\alpha,\beta,\eta,t)\coloneqq\frac{(1-\alpha)(1+\alpha+\beta)-(1-\alpha+\beta)\eta}{(1-\alpha+\beta)(1-\eta+\alpha+2\beta)}e^{-i(2+2\beta-\eta)t}+\frac{\beta(1+\alpha+\beta)-(1-\alpha+\beta)\eta}{(1-\alpha+\beta)(2\alpha+\beta-\eta)}e^{i(\eta-2\alpha)t}+\frac{(1+\alpha-\eta)\eta-\beta(1-\alpha-\eta)}{(2\alpha+\beta-\eta)(1+\alpha+2\beta-\eta)}e^{i(2\alpha+2\beta-\eta)t}$;
 $\mathrm{rhs}(x,y,z)\coloneqq e^{i(x-y-z)}-e^{i(y-x-z)}+e^{i(z-x-y)}$;
 if $\tau=\tau_{+}$ then
    $x^{\prime},\,y^{\prime},\,z^{\prime}\leftarrow x+c\tau,\,y+c\tau,\,c\tau-z$;
    $\eta\leftarrow(a-b)/(a+c);\quad t\leftarrow(a+c)\tau$;
    $(\alpha,\beta)\leftarrow\textsc{NSolve}(\mathrm{lhs}(\alpha,\beta,\eta,t)=\mathrm{rhs}(x^{\prime},y^{\prime},z^{\prime}))$ where $(\alpha,\beta)\in[0,1]\times[0,\infty)$ and $\alpha+\beta\geq\eta$;
    $\Omega_{1}\leftarrow 0$; $\Omega_{2}\leftarrow(a+c)\sqrt{(1-\alpha)\beta(1-\eta+\alpha+\beta)}$;
    $\delta\leftarrow-(a+c)\sqrt{\alpha(1+\beta)(\alpha+\beta-\eta)}$;
 else
    $x^{\prime},\,y^{\prime},\,z^{\prime}\leftarrow x-c\tau,\,y-c\tau,\,z-c\tau$;
    $\eta\leftarrow(a-b)/(a-c);\quad t\leftarrow(a-c)\tau$;
    $(\alpha,\beta)\leftarrow\textsc{NSolve}(\mathrm{lhs}(\alpha,\beta,\eta,t)=\mathrm{rhs}(x^{\prime},y^{\prime},z^{\prime}))$ where $(\alpha,\beta)\in[0,1]\times[0,\infty)$ and $\alpha+\beta\geq\eta$;
    $\Omega_{1}\leftarrow(a-c)\sqrt{(1-\alpha)\beta(1-\eta+\alpha+\beta)}$; $\Omega_{2}\leftarrow 0$;
    $\delta=(a-c)\sqrt{\alpha(1+\beta)(\alpha+\beta-\eta)}$;
   end if
end if
$H^{\prime\prime}_{1},\,H^{\prime\prime}_{2}\leftarrow(\Omega_{1}+\Omega_{2})X+\delta Z,\,(\Omega_{1}-\Omega_{2})X+\delta Z$;
@@
$H_{1},\,H_{2}\leftarrow U_{1}H^{\prime\prime}_{1}U_{1}^{\dagger}-H^{\prime}_{1},\,U_{2}H^{\prime\prime}_{2}U_{2}^{\dagger}-H^{\prime}_{2}$;
$A_{1},\,A_{2}\leftarrow A^{\prime}_{1}(A^{\prime\prime}_{1})^{\dagger}U_{1}^{\dagger},A^{\prime}_{2}(A^{\prime\prime}_{2})^{\dagger}U_{2}^{\dagger}$;
@@ 4.1. Protocol design (Hamiltonian description)
To illustrate the general applicability of our scheme, we consider the most general 2Q coupling Hamiltonian $H$ and local Rabi drives $H_{d_{i}}$ (Zhang and Whaley, 2005) for a pair of resonant qubits ($\omega_{q_{1}}=\omega_{q_{2}}=\omega$), with experimentally tunable pulse parameters $A_{i}(t),\delta$:
@@
where $(a,b,c)$ are the canonical coupling coefficients (Bennett et al., 2002) with $a\geq b\geq\lvert c\rvert$. We define the coupling strength
(3) $g:=a+b+\lvert c\rvert$
to facilitate the comparison of different physical platforms. For flux-tunable superconducting qubits, in the drive frame after the rotating wave approximation, $a=b=g/2$, $c=0$, $A_{i}(t)$ is the amplitude of a square wave envelope for a sinusoidal microwave drive, which we will assume is constant, and $\delta_{i}:=\frac{1}{2}(\omega_{d_{i}}-\omega)$ is the drive detuning (see (Krantz et al., 2019; Chen et al., 2024) for detailed derivations). We show that it is sufficient to set the drive frequencies to be identical, so $\delta_{1}=\delta_{2}=\delta$. For other physical platforms, $H_{d_{i}}$ may be realized via other means (e.g., lasers in Rydberg atoms or trapped ions). In this case, $A_{i}$ and $\delta_{i}$ may correspond to other physical parameters. Our framework can also be easily adapted to other possible $H_{d_{i}}$ (for example a $Y$ term). For mathematical convenience, we define $\Omega_{1,\,2}\coloneqq-\frac{1}{4}{(A_{1}\pm A_{2})}$, and we can rewrite the local Hamiltonians, which now act on the 2Q Hilbert space, as
(4) $H_{1}:=(\Omega_{1}+\Omega_{2})\,XI+\delta\,ZI,\,H_{2}:=(\Omega_{1}-\Omega_{2})\,IX+\delta\,IZ.$
@@
Algorithm 1 describes the details of our gate scheme. A target gate $U$ and the device 2Q Hamiltonian $H$ are provided as inputs. The goal is to determine the local Hamiltonians $H_{1}$ and $H_{2}$, with the interaction duration $\tau$, such that $e^{-i\tau(H+H_{1}+H_{2})}\sim U$. Specifically, the Weyl coordinates $(x,y,z)$ of $U$ are obtained via the KAK decomposition (line 1) according to Equation 1. The interaction Hamiltonian $H$, if not already in normal form, is transformed into its normal form (Dür et al., 2001) to obtain the coefficients $(a,b,c)$ (line 1):
$H=(U_{1}\otimes U_{2})(a\,XX+b\,YY+c\,ZZ)(U_{1}^{\dagger}\otimes U_{2}^{\dagger})+H^{\prime}_{1}+H^{\prime}_{2},$
where $U_{1},U_{2}$ are 1Q unitaries, $H_{1}^{\prime},H_{2}^{\prime}$ are 1Q Hermitian operators, and $a\geq b\geq|c|$. The parameters $(x,y,z)$ and $(a,b,c)$ are then fed to the subsequent solving procedure that computes $H_{1},H_{2},\tau$. With additional 1Q gate corrections $(A_{1},A_{2})$ and $(B_{1},B_{2})$ (line 1 and 1), finally
(5) $(A_{1}\otimes A_{2})\,e^{-i\tau(H+H_{1}+H_{2})}(B_{1}\otimes B_{2})$
exactly implements the desired $U$.
Herein we give a high-level explanation of Algorithm 1 while providing the detailed proof in Appendix A. First, the interaction time $\tau$ is determined as the maximum value among $\{\tau_{0},\tau_{+},\tau_{-}\}$. Which of the three is largest dictates the subscheme used:
  1. (1)
    $\tau_{0}$ for no-detuning (ND, $\delta=0$).
  2. (2)
    $\tau_{+}$ for equal-amplitude with opposite signs (EA+, $\Omega_{1}=0$), i.e., the sinusoidal drive amplitudes applied on qubit 1 and qubit 2 have equal magnitude but opposite signs.
  3. (3)
    $\tau_{-}$ for equal-amplitude with same sign (EA-, $\Omega_{2}=0$).
The three subschemes correspond to three distinct regions that partition the Weyl chamber (Chen et al., 2024). This partitioning of the Weyl chamber is convenient in terms of physical implementation but not unique.
The proof of the correctness of Algorithm 1 starts with transforming into the Bell (or “magic”) basis, where the canonical coupling Hamiltonian $aXX+bYY+cZZ$ becomes diagonal. In this basis, applying local drives effectively mixes different eigenstates. As explained in Appendix A, using the Bell basis motivates analyzing the unitary
(6) $V\coloneqq e^{-i\tau(H+H_{1}+H_{2})}\cdot YY$
whose eigenvalues are related to the Weyl coordinates of the gate that our gate scheme realizes. We want these eigenvalues to match the desired Weyl coordinates $(x,y,z)$ of $U$. Specifically, the “ND” subscheme leads to a special class of Hamiltonian for which $V$ has a simple analytic form. In contrast, the “EA+/EA-” subschemes involve equations whose solutions do not have a simple analytic form, as written in Algorithm 1. Furthermore, since the chosen duration $\tau$ matches the lower bound in (Hammerer et al., 2002), our gate scheme is time-optimal.
4.2. Numerical solving
Figure 4. $(\alpha,\,\beta)$ solution profiling for $\mathrm{SWAP}$ gate implementation under $\mathrm{XX}$ coupling ($a\neq 0,\,b=c=0$). All intersection points of orange and purple lines are valid solutions to the transcendental equations (“EA+” subscheme). Our numerical solver finds the optimal solution for hardware implementation (red point in the rightmost subfigure).
Figure 5. Compile-time singularity resolution via gate mirroring. (a) Gates in the near-identity region ($\mathcal{L}^{1}\leq r$) are mirrored to the $\mathrm{SWAP}$ corner. (b) Application to qft_4: In this example, $g_{i}\sim\mathrm{Can}(\pi/{2^{2+i}},0,0)$, such that $g_{2}$ and $g_{3}$ are assumed to be close enough to identity and are replaced by their mirrors while intermediate mappings are updated. Ultimately this resolution requires only one final mapping update $\pi_{f}$ and introduces no extra 2Q gate.
Each subscheme involves transcendental equations that lack analytical solutions and must be solved numerically. Furthermore, since $\mathrm{sinc}(x)$ function (lines 1 and 1) is not bijective, $S_{1}$ and $S_{2}$ have multiple solutions. To maximize the fidelity of a hardware implementation, we select by default the smallest solutions so that $\max\{|A_{1}|,|A_{2}|\}$ is minimized. While determining the primary roots for $S_{1}$ and $S_{2}$ is straightforward, the numerical solver for the EA+/EA-subschemes requires a more sophisticated approach. We employ a hybrid numerical optimization strategy that integrates a coarse grid search with a two-stage local refinement to solve for $(\alpha,\,\beta)$: ① Grid search initialization: We iterate over a predefined grid of initial guesses for $\alpha$ and $\beta$ both linearly spaced to adequately sample the parameter space. ② Two-stage refinement: For each $(\alpha,\,\beta)$ candidate in the grid search, we utilize the SLSQP method via scipy.optimize.minimize to minimize the residual between the expressions $\mathrm{lhs}$” (line 1) and $\mathrm{rhs}$” (line 1) but is of limited precision. The result is then refined to pinpoint the exact root using scipy.optimize.fsolve. ③ Optimal selection: Among the valid solutions, we select $(\alpha,\,\beta)$ that minimizes a “physical implementation penalty”, defined as the sum of the absolute values of the resulting pulse amplitudes $\Omega_{1}$ or $\Omega_{2}$ and detuning $\delta$.
This solving strategy achieves exceptional numerical precision and robustness. The Weyl coordinates computed from the returned solution has a mean absolute error on the order of $10^{-16}$ for ND and $10^{-13}$ for EA+/EA-. The resulting unitary produced has an infidelity on the order of $10^{-15}$, effectively reaching the limit of machine precision. In our field test with typical coupling Hamiltonians and common 2Q gates, as well as millions of random coupling Hamiltonians and target unitaries, the solving procedure consistently converges to high-precision solutions. Figure 4 illustrates the robustness of this procedure in identifying the primary $(\alpha,\beta)$ solution that corresponds to minimal pulse amplitudes and detuning for the SWAP gate under XX coupling.